# Collaborative filtering based on nonnegative/binary matrix factorization

### v3

@@ Abstract
Collaborative filtering generates recommendations by exploiting user–item similarities based on rating data, which often contains numerous unrated items. This paper proposes a nonnegative/binary matrix factorization (NBMF) algorithm modified for collaborative filtering and demonstrates that utilizing a low-latency Ising machine in NBMF is advantageous in terms of computation time. While previous studies have primarily applied NBMF to dense data, such as images, this study applies a modified NBMF to sparse data. Results show the benefits of using a low-latency Ising machine to implement the proposed method.
@@ II Problem formulation
where $W$ is a nonnegative real matrix. While $H$ is a binary matrix in which each element is 00 or $1$ for NBMF, it is a nonnegative real matrix for NMF. We assume that $n>k$ and $m>k$, which implies that NBMF and NMF provide low-rank matrix approximations of $V$. The rank constraint is helpful to prevent overfitting. Moreover, NBMF can be more resilient to overfitting due to the binary nature of matrix $H$.
In the context of collaborative filtering, $V$ is a rating matrix, where the $(i,j)$ element $v_{ij}$ represents user $i$’s rating of item $j$. In matrix $W$, each row corresponds to a user, while each column represents a basis vector associated with user preferences. In other words, $W$ consists of $k$ basis vectors, with each dimension being $n$. Meanwhile, each column of $H$ represents the coefficient vector related to the corresponding item. In NBMF, this coefficient vector indicates the combination of the selected basis vectors for the corresponding item. In general, the rating matrix contains numerous unrated entries. Matrix factorization techniques optimize $W$ and $H$ so that each rated entry in $V$ is well approximated by the corresponding element of $WH$. Then, each unrated entry in $V$ is estimated by the corresponding element of $WH$.
@@ III.1 Algorithm
The approach to conducting matrix factorization involves minimizing $\|V-WH\|_{F}$, where the Frobenius norm is defined as $\|A\|_{F}=\sqrt{\sum_{i,j}A_{ij}^{2}}$, and $A_{ij}$ is the $(i,j)$ element of $A$. To achieve minimization, NBMF employs an iterative alternative update procedure as follows:
@@ III.6 Data preparation
The method for extracting data at a specified filling rate is as follows. First, we sorted the columns in descending order by the percentage of filled elements in each column and then sorted the rows similarly. Next, we selected an $n\times m$ matrix whose $(1,1)$ element coincides with the first-row and first-column element of the sorted table, and calculated the filling rate of the matrix. By shifting the $(1,1)$-element location by one row and one column in the sorted table, we repeated the calculation of the filling rate. The $n\times m$ matrix with the closest filling rate to the desired filling rate was selected as the rating matrix.

### v4

@@ Abstract
Collaborative filtering generates recommendations by exploiting user–item similarities based on rating data, which often contains numerous unrated items. To predict scores for unrated items, matrix factorization techniques such as nonnegative matrix factorization (NMF) are often employed. Nonnegative/binary matrix factorization (NBMF), which is an extension of NMF, approximates a nonnegative matrix as the product of nonnegative and binary matrices. While previous studies have applied NBMF primarily to dense data such as images, this paper proposes a modified NBMF algorithm tailored for collaborative filtering with sparse data. In the modified method, unrated entries in the rating matrix are masked, enhancing prediction accuracy. Furthermore, utilizing a low-latency Ising machine in NBMF is advantageous in terms of the computation time, making the proposed method beneficial.
@@ II Problem formulation
where $W$ is a nonnegative real matrix. While $H$ is a binary matrix in which each element is 0 or 111 for NBMF, it is a nonnegative real matrix for NMF. We assume that $n>k$ and $m>k$, which implies that NBMF and NMF provide low-rank matrix approximations of $V$. The rank constraint is helpful to prevent overfitting. Moreover, NBMF can be more resilient to overfitting due to the binary nature of matrix $H$.
[…]
@@ III.1 Algorithm
[…]
@@ III.6 Data preparation
The method for extracting data at a specified filling rate is as follows. First, we sorted the columns in descending order by the percentage of filled elements in each column and then sorted the rows similarly. Next, we selected an $n\times m$ matrix whose (1,1)(1,1)( 1 , 1 ) element coincides with the first-row and first-column element of the sorted table, and calculated the filling rate of the matrix. By shifting the (1,1)(1,1)( 1 , 1 )-element location by one row and one column in the sorted table, we repeated the calculation of the filling rate. The $n\times m$ matrix with the closest filling rate to the desired filling rate was selected as the rating matrix.